Part A What is the numerical value of Kap if the equilibrium mixture has [Cu2+] and [CO, Express your answer using two significant figures equal to 12x10-5 M2 | ??? Submit Request Answer

media%2Fa8a%2Fa8a5c222-0f31-4e00-8fa7-48

media%2F025%2F0253aef9-5c85-4277-9bf7-f2

media%2F36f%2F36fec140-8770-4d50-ab2e-39
Answer #1

Part-A

[Cu+2]= [CO3-2]= 1.2x10^-5M

kSp = [Cu+2][CO3-2]

Ksp = (1.2x10^-5)^2

Ksp = 1.44x10^-10

part-B

PCl5(g) ---------------- PCl3(g) + Cl2(g)

0.65                           0               0

-x                               +x             +x

0.65-x                      +x                +x

according to the data

at equilibrium

number of moles of PCl3 = x= 0.17 mole

x= 0.17 moles

at equilibrium number moles of PCl5= 0.65-x= 0.65-0.17=0.48 moles

volume = 1.0L

[PCl5] = 0.48/1.0= 0.48M

[PCl5] = 0.48M

pRT-c

AT equlibrium

number of moles of Cl2= x= 0.17 moles

volume= 1.0L

[Cl2] = 0.17/1.0= 0.17M

part-D

Kc= [PCl3][Cl2]/[PCl5]

Kc= 0.17x0.17/0.48 =0.0602

Kc= 0.0602